Find the sum-of-products expansions of these boolean functions: F(x,y,z) = x.


1
Expert's answer
2022-05-24T18:19:45-0400

We should create the table for the function and take the Boolean sum of the minterm where the function evaluates to 1.

"F(x,y,z)=x"




So, the disjunctive normal form is "x\=y\=z+x\=yz+xy\=z+xyz"